Solve each system.
Number One:
y = 2x + 8
y = 3x - 1

Number Two:
2x - y = 2
2x - 2y = 4

Number Three:
-x + y = 2
2x + y = -1

Answer by stanbon(75887)

Number One:
y = 2x + 8
y = 3x - 1
-------------------
Substitute and solve for "x":
2x+8 = 3x-1
x = 9
---
Substitute to solve for "y":
y = 2*9=8
y = 26
===================================
Number Two:
2x - y = 2
2x - 2y = 4
--------
Subtract and solve for "y":
y = -2
Substitute to solve for "x":
2x + 2 = 2
2x = 0
x = 0
===================================
Number Three:
-x + y = 2
2x + y = -1
---
Subtract and solve for "x":
3x = -3
x = -1
---
Substitute to solve for "y":
1 + y = 2
y = 1
=====================================
Cheers,
Stan H.